# Machine Unlearning in Hyperbolic vs. Euclidean Multimodal Contrastive Learning: Adapting Alignment Calibration to MERU

### v1

@@ Abstract
Machine unlearning methods have become increasingly important for selective concept removal in large pre-trained models. While recent work has explored unlearning in Euclidean contrastive vision-language models, the effectiveness of concept removal in hyperbolic spaces remains unexplored. This paper investigates machine unlearning in hyperbolic contrastive learning by adapting Alignment Calibration to MERU, a model that embeds images and text in hyperbolic space to better capture semantic hierarchies. Through systematic experiments and ablation studies, we demonstrate that hyperbolic geometry offers distinct advantages for concept removal, achieving near perfect forgetting with reasonable performance on retained concepts, particularly when scaling to multiple concept removal. Our approach introduces hyperbolic-specific components including entailment calibration and norm regularization that leverage the unique properties of hyperbolic space. Comparative analysis with Euclidean models reveals fundamental differences in unlearning dynamics, with hyperbolic unlearning reorganizing the semantic hierarchy while Euclidean approaches merely disconnect cross-modal associations. These findings not only advance machine unlearning techniques but also provide insights into the geometric properties that influence concept representation and removal in multimodal models. Source code available at https://github.com/alex-pv01/HAC.
@@ 1 Introduction
It is commonly said that ’we learn through contrast’. By comparing seemingly opposite concepts, we gain deeper understanding—a fundamental principle underlying contrastive learning algorithms. This approach, which emphasizes identifying similarities and distinctions between data points, has emerged as a powerful framework for learning representations that effectively distinguish similarity from dissimilarity in machine learning systems [4].
The success of contrastive learning has led to the development of powerful multimodal models like CLIP, which aligns visual and textual representations in a shared Euclidean embedding space [22]. These models serve as foundational components for numerous downstream applications including zero-shot classification [21], image generation [24], and robotics [23]. More recently, MERU [7] has extended this paradigm by embedding images and text in hyperbolic rather than Euclidean space, better capturing the hierarchical relationships inherent in visual-semantic data.
Despite their impressive capabilities, these models rely on massive datasets scraped from the internet, which often contain problematic content including private information, copyrighted material, and harmful stereotypes [1]. This raises significant ethical and legal concerns, particularly as these models become more widely deployed. Machine unlearning—the ability to selectively remove information from trained models—initially emerged in response to regulatory frameworks like GDPR’s ”right to be forgotten”[8], but has since expanded to address a broader spectrum of challenges including privacy protection, safety enhancement, security concerns, fairness improvement, model robustness and copyright matters [28, 5].
In particular, concept removal (of NSFW or copyrighted content) becomes increasingly relevant to ensure a safe deployment for users and customers [20]. While recent work has explored unlearning in Euclidean contrastive learning models like CLIP [20, 29, 27, 14, 15], the unique geometry of hyperbolic spaces presents new challenges and opportunities that remain unexplored. The geometric distinction raises an important question: How does the underlying geometry affect the efficacy and dynamics of machine unlearning in contrastive learning models?
In this paper, we investigate the adaptation of Alignment Calibration (AC) to hyperbolic representations in MERU. We present a systematic comparison of unlearning effectiveness between CLIP (Euclidean) and MERU (hyperbolic) models trained on identical data. Our contributions are threefold:
@@
    Proposing Hyperbolic Alignment Calibration (HAC), an extension of AC to operate in hyperbolic space by reformulating its objective functions using hyperbolic distance metrics and incorporating entailment constraints.
@@
    Based on zero-shot classification evaluation, we conduct a comprehensive comparative analysis of machine unlearning in Euclidean and hyperbolic contrastive learning spaces using AC and HAC, revealing that the hyperbolic method excels at removing the targeted concept at the cost of a decrease in performance on the retain set.
@@
    Through visualization techniques, we further complement our insights on how the underlying geometric structures affect unlearning dynamics. Indicating that HAC may profit from the exponential expansion of the hyperbolic space to achieve better unlearning.
This work bridges two important research areas: geometric representation learning and machine unlearning. By analyzing how CLIP and MERU, with different geometries in the latent space, differ under the influence of an unlearning process. Providing insights into how representation spaces encode and forget concepts. Our adaptation of an existing unlearning method to hyperbolic geometry advances our understanding of the relationship between geometric structure and concept removal, contributing to the theoretical foundations of machine unlearning for advanced vision language models.
@@ 2 Related work
Our research lies at the intersection of machine unlearning, contrastive learning models, and geometric representation learning. This section reviews prior work in these areas and highlights the research gap we aim to address.
2.1 Geometric Representation Learning
The choice of geometry for representation spaces has significant implications for the expressive power of learned embeddings. While most machine learning models operate in Euclidean space, recent work has explored alternative geometries, particularly for data with hierarchical structure.
Hyperbolic geometry, characterized by its negative curvature, provides a continuous analog of tree-like structures [19]. In hyperbolic space, the volume of a ball grows exponentially with its radius, allowing to embed hierarchical data with low distortion. This property is valuable for representing data with natural hierarchies, such as taxonomies, social networks, and visual-semantic relationships [19, 10]. These geometric insights have direct implications for multimodal contrastive learning, where the choice of embedding space significantly impacts how concepts are organized and related across modalities.
2.2 Multimodal Contrastive Learning
Contrastive learning emerged as a powerful approach for unsupervised representation learning, training models to distinguish between similar (positive) and dissimilar (negative) pairs of examples [4]. This approach is formalized through the InfoNCE loss function [25]. Given a batch of $N$ samples, where for every sample $z_{i}$, there is a positive pair $z^{\prime}_{i}\neq z_{i}$, then:
$L_{\text{InfoNCE}}=\frac{-1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z_{i},z^{\prime}_{i})/\tau)}{\sum_{k=1,\;k\neq i}^{N}\exp(\text{sim}(z_{i},z_{k})/\tau)}$ (1)
where $\text{sim}(u,v)$ denotes cosine similarity, and $\tau$ is a temperature parameter. CLIP [22] extended this approach to the multimodal domain by learning a joint embedding space for images and text. More recently, MERU [7] enhanced CLIP’s representational power by embedding images and text in hyperbolic and using the negative hyperbolic distance as a similarity metric. It achieved improved performance on tasks involving hierarchical relationships while maintaining CLIP’s zero-shot capabilities.
@@ 2.2 Multimodal Contrastive Learning
2.3 Machine Unlearning
Machine Unlearning (MU) selectively removes the influence of specific training data from a trained model. It was introduced by Cao and Yang [2], as an approach to make systems forget targeted information. While early research focused primarily on classification tasks [28], the rapid growth of generative AI has expanded the scope of MU to concept removal [5, 9, 16].
Particularly, several works on concept removal for contrastive vision-language models have recently emerged. Poppi et al. [20] addressed safety concerns by developing a method to diminish CLIP’s sensitivity to NSFW content. Their approach is based on fine-tuning on synthetic ”unsafe” data. Kravets and Namboodiri [14] applied Lipschitz regularization to achieve zero-shot class unlearning in CLIP. In subsequent work, Kravets and Namboodiri [15] demonstrated that class forgetting in CLIP can be accomplished without any visual data by adapting the shared vision-text space, making the forgetting process more efficient. Yang et al. [29] introduced CLIPErase through three modules: a Forgetting Module that disrupts associations in the forget set, a Retention Module that preserves performance on the retain set, and a Consistency Module that maintains alignment with the original model. Wang et al. [27] proposed Alignment Calibration, which explicitly considers the properties of contrastive learning and optimizes toward novel auditing metrics to verify unlearning.
@@ 2.3 Machine Unlearning
Despite advances in geometric representation learning, multimodal contrastive learning, and machine unlearning, no prior work has investigated how hyperbolic geometry affects concept removal in contrastive models. This gap is significant given hyperbolic space’s natural capacity for representing hierarchical relationships, which could fundamentally alter how concepts are forgotten in multimodal representations.
@@ 3 Method
This section details our approach to adapting a machine unlearning technique to a different geometric space. We first formalize the unlearning problem, and then present our adaptation of AC from Euclidean to hyperbolic space.
@@ 3.1 Machine Unlearning for Concept Removal
Given a pre-trained model with parameters $\theta$, a dataset $\mathcal{D}$ with image-text pairs, and a concept $c$ to be forgotten, we define $\mathcal{D}_{f}$ as the subset containing instances that refer to $c$ and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ as the retain set. The unlearning task is to find parameters $\theta^{\prime}$ such that:
@@
    The model no longer associates visual and textual representations of concept $c$. Formally, for $(x,t)\in\mathcal{D}_{f}$, $\text{sim}_{\theta^{\prime}}(f_{\text{img}}(x),f_{\text{txt}}(t))\ll\text{sim}_{\theta}(f_{\text{img}}(x),f_{\text{txt}}(t))$.
@@
    Performance on other concepts remains unaffected. For $(x,t)\in\mathcal{D}_{r}$, $\text{sim}_{\theta^{\prime}}(f_{\text{img}}(x),f_{\text{txt}}(t))\approx\text{sim}_{\theta}(f_{\text{img}}(x),f_{\text{txt}}(t))$.
In this paper, we focus on concepts that correspond to discrete classes in common image-text datasets, call it concept-class. However, the formulation applies to more abstract concepts as well, provided that there is a proper way to obtain $D_{f}$.
@@ 3.2 Alignment Calibration in Euclidean Space
AC consists of a retaining loss and an unlearning loss that work together to selectively push apart image and text embeddings to forget, while preserving model performance [27]. Given a batch $\{(x^{r}_{i},t^{r}_{i})\}_{i=1}^{N}\subset\mathcal{D}_{r}$ and $\{(x^{f}_{i},t^{f}_{i})\}_{i=1}^{N^{\prime}}\subset\mathcal{D}_{f}$ from a dataset $\mathcal{D}$, for simplicity we take $N^{\prime}=N$, and $(x_{j},t_{i})\in\mathcal{D}$. The retention loss is taken from 15 and preserves model performance on $\mathcal{D}_{r}$:
@@
$L_{\text{neg}}=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\big{(}\text{sim}(x^{{}^{\prime}f}_{i},t^{{}^{\prime}f}_{j})+\text{sim}(x^{{}^{\prime}f}_{j},t^{{}^{\prime}f}_{i})\big{)}/\tau.$ (4)
This negative alignment calibration maximizes similarity between non-matching pairs in the forget set, confusing the model’s understanding of the concept.
@@
The positive alignment calibration minimizes similarity between matching pairs in the forget set, directly attacking the concept association.
@@
The performance preserving term ensures that general model capabilities remain intact. Note that this corresponds to the denominator from the original loss function 15
@@
$L_{\text{forget}}=-\alpha\cdot L_{\text{neg}}+\beta\cdot L_{\text{pos}}+\gamma\cdot L_{\text{perf}}.$ (8)
@@ 3.3 Alignment Calibration to Hyperbolic Space
HAC extends the AC approach to hyperbolic geometry by adapting each component of the loss function to operate in the Lorentz model of hyperbolic space. As in MERU’s formulation, we replace the cosine similarity with negative hyperbolic distance from component in AC. Additionally, HAC incorporates entailment terms to preserve and disrupt hierarchical relationships:
$L_{\text{r-ent}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\text{ext}(x^{{}^{\prime}r}_{i},t^{{}^{\prime}r}_{i})-\text{aper}(x^{{}^{\prime}r}_{i})),$ (10)
$L_{\text{f-ent}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\text{aper}(x^{{}^{\prime}f}_{i})-\text{ext}(x^{{}^{\prime}f}_{i},t^{{}^{\prime}f}_{i})),$ (11)
where $\text{ext}(x,y)$ is the exterior angle between the text embedding $x$, and the image embedding $y$, and $\text{aper}(x)$ is the half-aperture of the entailment cone for $x$. The retain-entailment loss preserves the hierarchical structure for concepts in the retain set, while the forget-entailment loss disrupts entailment relationships for concepts to be forgotten. The combined HAC loss is:
@@
Inspired by Li et al. [17] who shown that samples positioned close to the origin in hyperbolic space are likely to be out-of-distribution, we further propose adding a regularization term for the forgetting set that minimizes the hyperbolic norm of the image-text embeddings to be forgotten:
@@
where $||x||_{\mathcal{L}}=\sqrt{|\langle x,x\rangle_{\mathcal{L}}|}$ is the Lorentzian norm. This regularization pushes forget concepts to remain closer to the origin, additionally providing better numerical stability. The final HAC loss with regularization is:

### v2

@@ Abstract
Machine unlearning methods have become increasingly important for selective removal of harmful concepts in large multimodal models. While recent work has explored unlearning in Euclidean contrastive vision-language models, the effectiveness of concept removal in hyperbolic spaces remains unexplored. This paper explores this gap by adapting Alignment Calibration to MERU, a model that embeds images and text in hyperbolic space to better capture semantic hierarchies. Through systematic experiments and ablation studies, we demonstrate that hyperbolic geometry offers distinct advantages for concept removal, achieving lower accuracy at zero classification in forget set with reasonable performance on retain set, particularly when scaling to multiple concept removal. Our approach introduces hyperbolic-specific components including entailment calibration and norm regularization that leverage the unique properties of hyperbolic space. Comparative analysis with Euclidean models reveals fundamental differences in unlearning dynamics, with hyperbolic unlearning reorganizing the semantic hierarchy while Euclidean approaches merely disconnect cross-modal associations. These findings provide insights into the geometric properties that influence concept removal, setting a direction towards better control and safer deployment for large multimodal models. Code available at https://github.com/alex-pv01/HAC.
@@ 1 Introduction
It is commonly said that ’we learn through contrast’. By comparing seemingly opposite concepts, we gain deeper understanding—a fundamental principle underlying contrastive learning algorithms. This approach, has emerged as a powerful framework for learning representations that effectively distinguish similar from dissimilar in machine learning systems [6].
The success of contrastive learning has led to the development of powerful multimodal models like CLIP, which aligns visual and textual representations in a shared Euclidean embedding space [30]. These models serve as foundational components for numerous downstream applications including zero-shot classification [29], image generation [32], and robotics [31]. More recently, MERU [9] has extended this paradigm by embedding images and text in hyperbolic space, better capturing the hierarchical relationships inherent in visual-semantic data.
Despite their impressive capabilities, these models rely on massive datasets scraped from the internet, which often contain problematic content including private information, copyrighted material, and harmful data [2]. This raises significant ethical and legal concerns, as these models become widely deployed. Machine unlearning—the ability to selectively remove information from trained models—initially emerged in response to regulatory frameworks like GDPR’s ”right to be forgotten”[10], but has since expanded to address a broader spectrum of challenges including privacy protection, safety enhancement, security concerns, fairness improvement, model robustness and copyright matters [37, 7].
In particular, concept removal (of NSFW or copyrighted content) has been increasingly relevant to ensure a safe deployment [28]. While recent work has explored concept removal in Euclidean contrastive learning models like CLIP [28, 38, 36, 16, 17], the unique geometry of hyperbolic spaces presents new challenges and opportunities that remain unexplored. The geometric distinction raises an important question: How does the underlying geometry affect the efficacy and dynamics of machine unlearning in contrastive learning models?
In this paper, we investigate the adaptation of Alignment Calibration (AC) to hyperbolic representations in MERU. We present a systematic comparison of unlearning effectiveness between CLIP and MERU models trained on identical data. Our contributions are threefold:
@@
    Proposing Hyperbolic Alignment Calibration (HAC), an extension of AC to operate in hyperbolic space by reformulating its objective functions using hyperbolic distance and incorporating entailment constraints.
@@
    Based on zero-shot classification evaluation, we conduct a comparative analysis of machine unlearning in Euclidean and hyperbolic contrastive learning spaces using AC and HAC, revealing that the hyperbolic method excels at removing the targeted concept at the cost of a decrease in performance on the retain set.
@@
    Through visualization techniques, we complement our insights on how the geometries affect unlearning dynamics. Indicating that HAC may profit from the exponential expansion of the hyperbolic space to achieve better unlearning.
This work bridges two important research areas: geometric representation learning and machine unlearning. Our study advances the understanding on how representation spaces encode and forget concepts, contributing to the theoretical foundations of machine unlearning for advanced vision language models.
@@ 2 Related work
Hyperbolic Representation Learning
The choice of geometry for representation spaces has significant implications for the expressive power of learned embeddings. While most machine learning models operate in Euclidean space, recent work has explored alternative geometries, particularly hyperbolic geometry for data with hierarchical structure [27, 23].
Characterized by its negative curvature, a hyperbolic space provides a continuous analog of tree-like structures [24]. In hyperbolic space, the volume of a ball grows exponentially with its radius, allowing to embed hierarchical data with low distortion [33]. This property is valuable for representing data with natural hierarchies, such as taxonomies, social networks, and visual-semantic relationships [24, 12, 9]. These geometric insights have direct implications for multimodal contrastive learning, on how concepts are organized and related across modalities.
Multimodal Contrastive Learning
Contrastive learning emerged as a powerful approach for unsupervised representation learning, training models to distinguish between similar (positive) and dissimilar (negative) pairs of examples [6]. This approach is formalized through the InfoNCE loss function [34]. Given a batch of $N$ samples, where for every sample $z_{i}$, there is a positive pair $z_{k_{i}}\neq z_{i}$, then:
$L_{\text{InfoNCE}}=\frac{-1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z_{i},z_{k_{i}})/\tau)}{\sum_{k=1,\;k\neq i}^{N}\exp(\text{sim}(z_{i},z_{k})/\tau)}$ (1)
where $\text{sim}(u,v)$ denotes cosine similarity, and $\tau$ is a temperature parameter. CLIP [30] extended this approach to the multimodal domain by learning a dual-encoder architecture with separate visual and textual encoders that project images and their corresponding text descriptions into a shared Euclidean space. More recently, MERU [9] enhanced CLIP’s representational power by embedding images and text in hyperbolic and using the negative hyperbolic distance as a similarity metric. It achieved improved performance on tasks involving hierarchical relationships while maintaining CLIP’s zero-shot capabilities.
@@
Machine Unlearning
Machine Unlearning (MU), introduced by Cao and Yang [4], selectively removes the influence of specific training data from a trained model. While early research focused primarily on classification tasks [37], the rapid growth of generative AI has expanded the scope of MU to higher-order or latent information [7, 11, 20].
Particularly, several works on concept removal for contrastive vision-language models have recently emerged. Poppi et al. [28] addressed safety concerns by developing a method to diminish CLIP’s sensitivity to NSFW content. Their approach is based on redirecting synthetic ”unsafe” data to match ”safe” data. Kravets and Namboodiri [16] applied Lipschitz regularization to achieve zero-shot class unlearning in CLIP. In subsequent work, Kravets and Namboodiri [17] demonstrated that class forgetting in CLIP can be accomplished without any visual data by adapting the shared vision-text space, making the forgetting process more efficient. Yang et al. [38] introduced CLIPErase through three modules: a Forgetting Module that disrupts associations in the forget set, a Retention Module that preserves performance on the retain set, and a Consistency Module that maintains alignment with the original model. Similarly, Wang et al. [36] proposed Alignment Calibration, which explicitly considers the properties of contrastive learning and introduce toward novel auditing metrics to verify unlearning.
@@
Despite mentioned, no prior work has investigated how hyperbolic geometry affects concept removal in contrastive models. This gap is significant given hyperbolic space’s natural capacity for representing hierarchical relationships, which could fundamentally alter how concepts are forgotten in multimodal representations.
@@ 3.1 Machine Unlearning for Concept Removal
Given a pre-trained contrastive model with image encoder $f_{\text{img}}$, text encoder $f_{\text{txt}}$, and a training dataset $\mathcal{D}$ with image-text pairs. Let $c$ be the concept to be removed, we define $\mathcal{D}_{f}$ as the subset containing instances that refer to $c$ and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ as the retain set. That is, if $c$ is the concept of ”dog”, $D_{f}$ consists of positive pairs of images of dogs and captions describing the image. The unlearning task is to find modify the original model and obtain $f^{*}_{\text{img}},f^{*}_{\text{txt}}$ such that:
@@
    The model no longer associates visual and textual representations of concept $c$. Formally, for a positive pair $(x_{f},t_{f})\in\mathcal{D}_{f}$ exists a pair $(x,t)\in\mathcal{D}_{r}$,
    $\text{sim}(f^{*}_{\text{img}}(x_{f}),f^{*}_{\text{txt}}(t_{f}))\ll\text{sim}(f^{*}_{\text{img}}(x_{f}),f^{*}_{\text{txt}}(t)),$ (2)
    and
    $\text{sim}(f^{*}_{\text{img}}(x_{f}),f^{*}_{\text{txt}}(t_{f}))\ll\text{sim}(f^{*}_{\text{img}}(x),f^{*}_{\text{txt}}(t_{f})).$ (3)
@@
    Performance on other concepts in remains unaffected. For positive pairs $(x,t)\in\mathcal{D}_{r}$,
    $\text{sim}(f^{*}_{\text{img}}(x),f^{*}_{\text{txt}}(t))\approx\text{sim}(f_{\text{img}}(x),f_{\text{txt}}(t)).$ (4)
Note that concept removal is a recent problem highlighted in literature and has no established formal definition [7]. However, the above definition aligns with contrastive learning approaches on related work [36, 16, 17], where unlearning is measured by accuracy on zero-shot classification, as later discussed in Section 4.2.
@@ 3.2 Alignment Calibration in Euclidean Space
AC consists of a retaining loss and an unlearning loss that work together to selectively push apart image and text embeddings to forget, while preserving model performance [36]. Given a batch $\{(x^{r}_{i},t^{r}_{i})\}_{i=1}^{N}\subset\mathcal{D}_{r}$ and $\{(x^{f}_{i},t^{f}_{i})\}_{i=1}^{N}\subset\mathcal{D}_{f}$ from a dataset $\mathcal{D}$. Let $x^{\prime}=f_{\text{img}}(x)$ and $t^{\prime}=f_{\text{txt}}(t)$ denote image and text embeddings respectively. The retention loss consists of the usual loss for CLIP (see Sec. 10.1 in supp. material), preserving model performance on $\mathcal{D}_{r}$:
@@
  1. 1.
    Negative alignment calibration (Eq. 7), maximizes similarity between negative pairs in the forget set, confusing the model’s understanding of $c$.
    $L_{\text{neg}}=-\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\frac{\text{sim}(x^{{}^{\prime}f}_{i},t^{{}^{\prime}f}_{j})+\text{sim}(x^{{}^{\prime}f}_{j},t^{{}^{\prime}f}_{i})}{\tau}.$ (7)
  2. 2.
    Positive alignment calibration (Eq. 8), minimizes similarity between matching pairs in the forget set, directly attacking the concept association.
@@
  3. 3.
    Performance preserving (Eq. 9), ensures that general model capabilities remain intact. Note that this corresponds to the denominator from the retention loss, Eq. 5, but also considering $D_{f}$.
@@
$L_{\text{forget}}=\alpha\cdot L_{\text{neg}}+\beta\cdot L_{\text{pos}}+\gamma\cdot L_{\text{perf}}.$ (11)
@@ 3.3 Alignment Calibration to Hyperbolic Space
HAC extends the AC approach to hyperbolic geometry by adapting each component of the loss function to operate in the Lorentz model of hyperbolic space. As in MERU’s formulation (see Sec. 10.2 in supp. material), we replace the cosine similarity with negative hyperbolic distance from component in AC. HAC incorporates entailment terms, introduced by Desai et al. [9]. We propose adapting the entailment to preserve and disrupt hierarchical relationships:
$L_{\text{r-ent}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\text{ext}(x^{{}^{\prime}r}_{i},t^{{}^{\prime}r}_{i})-\text{aper}(t^{{}^{\prime}r}_{i})),$ (13)
$L_{\text{f-ent}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\text{aper}(t^{{}^{\prime}f}_{i})-\text{ext}(x^{{}^{\prime}f}_{i},t^{{}^{\prime}f}_{i})),$ (14)
where $\text{ext}(x,t)$ is the exterior angle between the text embedding $t$, and the image embedding $x$, and $\text{aper}(t)$ is the half-aperture of the entailment cone for $t$. The retain-entailment loss (Eq. 13) preserves the hierarchical structure for concepts in the retain set, minimizing the angle difference when the embedded image $x^{{}^{\prime}r}_{i}$ lies outside the entailment cone for the text $t^{{}^{\prime}r}_{i}$, following the same logic as MERU. On the other had, the forget-entailment loss (Eq. 14) disrupts entailment relationships for texts and images of the concepts to be forgotten, minimizing the angle difference when $x^{{}^{\prime}f}_{i}$ is inside the entailment cone of $t^{{}^{\prime}f}_{i}$, forcing the images to move close to the edge of the cone. The combined HAC loss is:
@@
Additionally, inspired by Li et al. [21], who show that samples positioned close to the origin in hyperbolic space are likely to be out-of-distribution, we further propose adding a regularization term for the forgetting set that minimizes the hyperbolic norm of the image-text embeddings to be forgotten:
@@
where $||\cdot||_{\mathcal{L}}$ is the Lorentzian norm, see Sec. 10.2. This regularization pulls instances to remain closer to the origin, which in turn provides better numerical stability [14]. The final HAC loss with regularization is: